# Complexity reduction of physical models: an equation-free approach by means of scaling

### v1

@@ 1 Introduction
When dealing with complex models, it is often useful to consider an asymptotic analysis with the aim of simplifying the model under consideration. Such an approach identifies a parameter $\varepsilon>0$ and investigates the limit for $\varepsilon\to 0$ of the solutions to the given equations. When $\varepsilon$ is small, the solutions to the model considered in first place can be accurately reproduced by the solutions of the same equations with $\varepsilon$ formally set to 00, i.e., with some terms neglected. However, in general, it is hard to understand how small such an $\varepsilon$ must be to consider it “small enough” and, then, to allow the use of solutions to the simplified model.
@@
In particular, for models involving very different scales or over-parametrization, reducing complexity and applying appropriate scaling becomes crucial for preventing instability in calibration [1, 2, 3]. This turns to be even more critical when feeding synthetic data from physical models into machine learning models [4], as imbalances or miscalibrations in the data can significantly affect model performance [5]. Proper scaling ensures that synthetic data reflects realistic parameter distributions, helping to mitigate overfitting and instability in parameter estimation [1]. Ultimately, scaling improves calibration, stabilizes predictions and may lead to more reliable machine learning model training, even when using synthetic data.
@@
Our approach is based on already introduced methodologies, that we called Optimal Scaling (OS) [6] and Optimal Scaling with Constraints (OSC) [7]. Similarly to traditional nondimensionalization approaches, those allow recasting dimensional equations into unitless form. However, in contrast to traditional techniques, they do not require a problem-dependent insight into the phenomenon under consideration. In addition, they ensure a strong reduction of ranges of magnitudes assumed by the equations’ coefficients, provide accurate estimation of characteristic features of considered systems and allow reducing round-off and numerical errors.
@@
Such approaches were applied in [6, 7] to the prospective Population Balance Model [8] for Latex Particles Morphology Formation (LPMF PBM) [9, 10]. First, in [6], we obtained dimensionless scaled equations with minimized variation of the relevant quantities. Such a feature was beneficial in diminishing numerical errors, and avoiding unphysical behavior of the computed solution. Second, in [7], a quantitative criterion for locating regions of slow and fast aggregation was introduced to derive a family of dimensionless LPMF PBM of reduced complexity. Moreover, in [6], the Optimal Scaling was also applied to liquid crystal and quantum models, helping to decrease the range of magnitudes assumed by the involved quantities, without compromising on accuracy of characteristic features of considered systems [11, 12].
@@
Here, we present an advanced variant of the Optimal Scaling approaches which inherits benefits of its predecessors and, in addition, takes control of over-parametrization. We refer to this method as Generalized Optimal Scaling (GOS). The proposed methodology can provide a quantitative criterion for identifying those values of physical parameters that allow discarding a few terms in the considered equations without significantly modifying their solutions. This technique does not depend on the model, and as such its application to any specific model will produce results which exhibit a certain dependency on the model, that needs to be investigated numerically. In order to concretely illustrate our derivations and findings, we consider the classical projectile model.
@@
The paper is structured as follows. Section 2 clarifies the employed notations and summarizes the projectile model we use in this work as a case study. Section 3 outlines the main achievements of the proposed scaling methods [6, 7] and puts the OSC method, previously formulated in the context of the LPMF PBM, into a general framework. Section 4 improves further the generalized OSC by extending its functionalities with ability to reduce to its minimum the number of parameters employed in a model. A detailed formulation of the resulting GOS method is presented in Table 1 in comparison with the original OSC method. In Section 5, we test the Optimal Scaling methodologies on the projectile model. Section 6 provides conclusions and discusses the capabilities of presented techniques.
@@ 2.1 Employed Notation
    being $[\tilde{q}]$ the dimensions of $\tilde{q}$, i.e., its units of measure, and $q$ the numerical value associated with $\tilde{q}$ in the chosen units $[\tilde{q}]$.
@@ 2.2 The classical projectile model: scaling strategies
A ball is thrown vertically upward from a certain height above the surface of the Earth. The height $\tilde{y}(\tilde{t})$ reached at time $\tilde{t}$ is measured in meters, while the elapsed time $\tilde{t}$ in seconds, i.e., $[\tilde{y}]=\mathrm{m}$, $[\tilde{t}]=\mathrm{s}$. As presented in [13], the variable $\tilde{y}$ satisfies the Ordinary Differential Equation
@@
being $\tilde{g}$ the gravitational acceleration, $\tilde{R}$ the radius of the Earth, $\tilde{h}_{0}$ and $\tilde{v}_{0}$ the initial height and velocity, respectively. As indicated in [14], the dimensionless counterpart of (2) is obtained by the change of variables
@@
Now it is possible to optimize, say, the coefficients $\lambda_{2},\lambda_{3},\lambda_{4}$ as functions of $\pi_{1},\pi_{2}$ (instead of $\tilde{\theta}_{1},\tilde{\theta}_{2}$, as suggested in [6, 7]). If one aims to have $\lambda_{2},\lambda_{3},\lambda_{4}$ as close as possible to $1$ (the approach taken by OS), it follows:
@@ 3 Optimal Scaling: Overview
where $\tilde{F}$ is a function (with dimensions $[\tilde{F}]$) of unknown variables and independent quantities $\tilde{x}_{1},\dots,\tilde{x}_{N_{\mathsf{x}}}$ and physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, whose numerical values
@@
In the literature [13, 14], it is often suggested to set $N_{\mathsf{x}}$ out of the $N_{\mathsf{l}}$ coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ equal to $1$. Then, one aims to solve the resulting system of equations with respect to the values $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}$ of the characteristic constants. This approach presents some deficiencies, such as, e.g., at most $N_{\mathsf{x}}$ out of $N_{\mathsf{l}}$ independent quantities can be enforced to the value of $1$, with no control on the rest of the dimensionless coefficients. Furthermore, the selection of coefficients enforced to the value of $1$ depends on the particular problem with no general guidelines given so far.
@@
In this paper, we address these issues by proposing techniques for selection of numerical values $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}\in(0,\infty)$ of the characteristic constants $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ to scale the dimensional model (9) effectively, according to the objectives set for the specific scaling method (see Table 1 from Section 4). We remark that such procedures do not require a physical insight into the considered phenomenon and corresponding equations. The main features and achievements of presented methodologies are summarized in what follows.

### v2

@@ 1 Introduction
When dealing with complex models, it is often useful to consider an asymptotic analysis with the aim of simplifying the model under consideration. Such an approach identifies a parameter $\varepsilon>0$ and investigates the limit for $\varepsilon\to 0$ of the solutions to the given equations. When $\varepsilon$ is small, the solutions to the model considered in first place can be accurately reproduced by the solutions of the same equations with $\varepsilon$ formally set to 0, i.e., with some terms neglected. However, in general, it is hard to understand how small such an $\varepsilon$ must be to consider it “small enough” and, then, to allow the use of solutions to the simplified model.
@@
In particular, for models involving very different scales or over-parametrization, reducing complexity and applying appropriate scaling becomes crucial for preventing instability in calibration [https://doi.org/10.1111/1467-9868.00294, https://doi.org/10.1038/s41467-021-26107-z, ZHANG2024109854]. This turns to be even more critical when feeding synthetic data from physical models into machine learning models [technologies9030052], as imbalances or miscalibrations in the data can significantly affect model performance [5128907]. Proper scaling ensures that synthetic data reflects realistic parameter distributions, helping to mitigate overfitting and instability in parameter estimation [https://doi.org/10.1111/1467-9868.00294]. Ultimately, scaling improves calibration, stabilizes predictions and may lead to more reliable machine learning model training, even when using synthetic data.
@@
Our approach is based on already introduced methodologies, that we called Optimal Scaling (OS) [RUSCONI2019106944] and Optimal Scaling with Constraints (OSC) [RUSCONI2023127756]. Similarly to traditional nondimensionalization approaches, those allow recasting dimensional equations into unitless form. However, in contrast to traditional techniques, they do not require a problem-dependent insight into the phenomenon under consideration. In addition, they ensure a strong reduction of ranges of magnitudes assumed by the equations’ coefficients, provide accurate estimation of characteristic features of considered systems and allow reducing round-off and numerical errors.
@@
Such approaches were applied in [RUSCONI2019106944, RUSCONI2023127756] to the prospective Population Balance Model [Ramkrishna2000] for Latex Particles Morphology Formation (LPMF PBM) [DDPM_2016, PhDThesis_Rusconi_PMCQS]. First, in [RUSCONI2019106944], we obtained dimensionless scaled equations with minimized variation of the relevant quantities. Such a feature was beneficial in diminishing numerical errors, and avoiding unphysical behavior of the computed solution. Second, in [RUSCONI2023127756], a quantitative criterion for locating regions of slow and fast aggregation was introduced to derive a family of dimensionless LPMF PBM of reduced complexity. Moreover, in [RUSCONI2019106944], the Optimal Scaling was also applied to liquid crystal and quantum models, helping to decrease the range of magnitudes assumed by the involved quantities, without compromising on accuracy of characteristic features of considered systems [Gartland2018, PhysRevA.53.2135].
@@
The purpose of this study is to generalize the already introduced methodologies and demonstrate in a transparent way the use and advantages of such an approach on the classical benchmark in the scaling theory. In particular, we present an advanced variant of the Optimal Scaling approaches which inherits benefits of its predecessors and, in addition, takes control of over-parametrization. We refer to this method as Generalized Optimal Scaling (GOS). The proposed methodology can provide a quantitative criterion for identifying those values of physical parameters that allow discarding a few terms in the considered equations without significantly modifying their solutions. This technique does not depend on the model, and as such its application to any specific model will produce results which exhibit a certain dependency on the model, that needs to be investigated numerically. In order to concretely illustrate our derivations and findings, we consider the classical projectile model.
@@
The paper is structured as follows. Section 2 clarifies the employed notations and summarizes the projectile model we use in this work as a case study. Section 3 outlines assumptions and statements formulated and proven in this study to support the presented scaling approaches. Moreover, Section 3 sums up the main achievements of the proposed methodologies [RUSCONI2019106944, RUSCONI2023127756] and puts the OSC method, previously formulated in the context of the LPMF PBM, into a general framework. Section 4 improves further the generalized OSC by extending its functionalities with ability to reduce to its minimum the number of parameters employed in a model. A detailed formulation of the resulting GOS method is presented in Table 1 in comparison with the original OSC method. In Section 5, we test the Optimal Scaling methodologies on the projectile model. Section 6 provides conclusions and discusses the capabilities of presented techniques.
@@ 2.1 Employed Notation
    being $\tilde{q}$] the dimensions of $\tilde{q}$, i.e., its units of measure, and $q$ the numerical value associated with $\tilde{q}$ in the chosen units $\tilde{q}$].
@@ 2.2 The classical projectile model: scaling strategies
[…]
@@
being $\tilde{g}$ the gravitational acceleration, $\tilde{R}$ the radius of the Earth, $\tilde{h}_{0}$ and $\tilde{v}_{0}$ the initial height and velocity, respectively. As indicated in [Scaling_Langtangen], the dimensionless counterpart of (2) is obtained by the change of variables
@@
Now it is possible to optimize, say, the coefficients $\lambda_{2},\lambda_{3},\lambda_{4}$ as functions of $\pi_{1},\pi_{2}$ (instead of $\tilde{\theta}_{1},\tilde{\theta}_{2}$, as suggested in [RUSCONI2019106944, RUSCONI2023127756]). If one aims to have $\lambda_{2},\lambda_{3},\lambda_{4}$ as close as possible to $1$ (the approach taken by OS), it follows:
@@ 3 Optimal Scaling: Overview
where $\tilde{F}$ is a function (with dimensions $\tilde{F}$]) of unknown variables and independent quantities $\tilde{x}_{1},\dots,\tilde{x}_{N_{\mathsf{x}}}$ and physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, whose numerical values
@@
In the literature [Holmes2009_ND, Scaling_Langtangen], it is often suggested to set $N_{\mathsf{x}}$ out of the $N_{\mathsf{l}}$ coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ equal to $1$. Then, one aims to solve the resulting system of equations with respect to the values $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}$ of the characteristic constants. This approach presents some deficiencies, such as, e.g., at most $N_{\mathsf{x}}$ out of $N_{\mathsf{l}}$ independent quantities can be enforced to the value of $1$, with no control on the rest of the dimensionless coefficients. Furthermore, the selection of coefficients enforced to the value of $1$ depends on the particular problem with no general guidelines given so far.
@@
In this paper, we address these issues by proposing techniques for selection of numerical values $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}\in(0,\infty)$ of the characteristic constants $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ to scale the dimensional model (9) effectively, according to the objectives set for the specific scaling method (see Table 1 from Section 4). We remark that such procedures do not require a physical insight into the considered phenomenon and corresponding equations.
3.1 Assumptions and Main Results
The scaling approaches presented in this study rely on the Assumptions listed below. Such assumptions give rise to the Theorems formulated in Section 3.1.1 to constitute the tools for the nondimensional setting of the dimensional model (9). In addition, the Theorems stated in Section 3.1.2 indicate how to optimize the coefficients of the resulting dimensionless model (12).
Assumption 1.
$\kappa_{i},\theta_{j}\in(0,\infty),\quad c_{i,j}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{l}},\quad\forall j=1,\dots,N_{\mathsf{x}},$ (14)
where each $\tilde{\kappa}_{i}$ is an explicit function of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ only, and all the exponents $c_{i,j}$ in (14) are explicitly known real numbers.
Assumption 2.
Each of the factors $\tilde{\kappa}_{1},\dots,\tilde{\kappa}_{N_{\mathsf{l}}}$ in (14) can be written as a power-law monomial of the physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ only:
$\tilde{\kappa}_{i}=\tilde{p}_{1}^{d_{i,1}}\,\tilde{p}_{2}^{d_{i,2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{d_{i,N_{\mathsf{p}}}},\quad d_{i,j}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{l}},\quad\forall j=1,\dots,N_{\mathsf{p}},$ (15)
with all the exponents $d_{i,j}$ in (15) to be explicitly known real numbers.
Assumption 3.
$\tilde{\theta}_{j}=\tilde{p}_{1}^{t_{1,j}}\,\tilde{p}_{2}^{t_{2,j}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{t_{N_{\mathsf{p}},j}},\quad t_{i,j}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{p}},\quad\forall j=1,\dots,N_{\mathsf{x}},$ (16)
where the exponents $t_{i,j}$ in (16) are unknowns of scaling approaches.
Assumption 4.
Assumption 5.
Assumption 6.
$\tilde{p}_{j}]=\mathrm{u}_{1}^{m_{1,j}}\,\mathrm{u}_{2}^{m_{2,j}}\,\cdots\,\mathrm{u}_{N_{\mathsf{u}}}^{m_{N_{\mathsf{u}},j}}\neq 1,\quad m_{i,j}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{u}},\quad\forall j=1,\dots,N_{\mathsf{p}},$ (17)
with $\mathrm{u}_{1}^{\beta_{1}}\,\mathrm{u}_{2}^{\beta_{2}}\,\cdots\,\mathrm{u}_{N_{\mathsf{u}}}^{\beta_{N_{\mathsf{u}}}}=1\Leftrightarrow\beta_{1}=\beta_{2}=\cdots=\beta_{N_{\mathsf{u}}}=0$, and all the exponents $m_{i,j}$ in (17) being explicitly known real numbers.
Aiming to minimize the number of involved quantities, one can also require the independence of the $N_{\mathsf{l}}$ coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13), or, equivalently, verify that the rank of the matrix made by the exponents in (14) and (15) is equal to $N_{\mathsf{l}}$, i.e.,
$\mathrm{rank}\left(\begin{pmatrix}c_{1,1}&\cdots&c_{1,N_{\mathsf{x}}}&d_{1,1}&\cdots&d_{1,N_{\mathsf{p}}}\\ c_{2,1}&\cdots&c_{2,N_{\mathsf{x}}}&d_{2,1}&\cdots&d_{2,N_{\mathsf{p}}}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ c_{N_{\mathsf{l}},1}&\cdots&c_{N_{\mathsf{l}},N_{\mathsf{x}}}&d_{N_{\mathsf{l}},1}&\cdots&d_{N_{\mathsf{l}},N_{\mathsf{p}}}\end{pmatrix}\right)=N_{\mathsf{l}}.$ (18)
In Section 5, we show that Assumptions˜1, 2, 4, 5 and 6 hold when considering the projectile model presented in Section 2.2. For such a physical model, we also make Assumption˜3, as explicitly indicated by (47). Though such an assumption can be motivated by its dimensional consistency, there exist physical models satisfying Assumptions˜1, 2, 4, 5 and 6, while violating Assumption˜3 (see, for example, the equation for energy density (2.49) in [Stewart2004]).
3.1.1 Nondimensional Setting
Theorem 1 algebraically formalizes a criterion for dimensionlessness of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$. The proof is shown in Appendix A.
Theorem 1.
Given Assumption 6 for physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, any quantity of the form
$\tilde{p}_{1}^{z_{1}}\,\tilde{p}_{2}^{z_{2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{z_{N_{\mathsf{p}}}},\quad\mbox{with}\quad z_{1},z_{2},\dots,z_{N_{\mathsf{p}}}\in\mathbb{R},$ (19)
is dimensionless if and only if the vector $\vec{z}\coloneqq(z_{1},z_{2},\dots,z_{N_{\mathsf{p}}})^{T}\in\mathbb{R}^{N_{\mathsf{p}}}$ of exponents in (19) belongs to the null space (kernel) of the matrix
$\textbf{M}\coloneqq\begin{pmatrix}m_{1,1}&m_{1,2}&\cdots&m_{1,N_{\mathsf{p}}}\\ m_{2,1}&m_{2,2}&\cdots&m_{2,N_{\mathsf{p}}}\\ \vdots&\vdots&\ddots&\vdots\\ m_{N_{\mathsf{u}},1}&m_{N_{\mathsf{u}},2}&\cdots&m_{N_{\mathsf{u}},N_{\mathsf{p}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{u}}\times N_{\mathsf{p}}},$ (20)
where the entries of M (20) are the exponents in (17) specified by Assumption 6.
Theorem 2 determines the minimal number of independent dimensionless parameters and indicates a way to compute them in terms of the provided physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$. The proof is shown in Appendix B.
Theorem 2.
Given Assumptions 4 to 6 for physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, there exist only $N_{\mathsf{0}}=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})$ independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$ of the form (19), with M as in (20). Such parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ are identified by the vectors composing a basis of the null space (kernel) of the matrix M (20):
$\mbox{with}\quad\vec{b}_{1},\vec{b}_{2},\dots,\vec{b}_{N_{\mathsf{0}}}\quad\mbox{constituing a}\quad\mathrm{basis}(\mathrm{ker}(\textbf{M})),$ (21)
Theorem 3 states that the dimensionless coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be written in terms of the independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (21) only. The proof is shown in Appendix C.
Theorem 3.
Given Assumptions 1 to 6 and $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ (21), each of the dimensionless coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13) can be written as a power-law monomial of the independent dimensionless parameters $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}$ only, i.e.,
$\lambda_{i}=\pi_{1}^{\alpha_{i,1}}\,\pi_{2}^{\alpha_{i,2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{\alpha_{i,N_{\mathsf{0}}}},\quad\alpha_{i,k}\in\mathbb{R},\quad\forall i=1,\dots,N_{\mathsf{l}},\quad\forall k=1,\dots,N_{\mathsf{0}},$ (22)
where the exponents $\alpha_{i,k}$ in (22) are unknowns of scaling procedures.
Theorem 4 offers a way to obtain the unknown exponents $t_{i,j}$ in (16). The proof is shown in Appendix D.
Theorem 4.
Given Assumptions 1 to 6, the vector
$\vec{t}\coloneqq\begin{pmatrix}\vec{t}_{1}\\ \vec{t}_{2}\\ \vdots\\ \vec{t}_{N_{\mathsf{x}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{p}}\,N_{\mathsf{x}}},\quad\mbox{with}\quad\vec{t}_{j}\coloneqq\begin{pmatrix}t_{1,j}\\ t_{2,j}\\ \vdots\\ t_{N_{\mathsf{p}},j}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{p}}},$ (23)
of unknown exponents in (16) must satisfy the linear system
$\textbf{Y}\,\vec{t}=\vec{\gamma},\quad\textbf{Y}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})},\quad\vec{\gamma}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{p}}},$ (24)
where the entries of Y are explicitly known real numbers, while the entries of $\vec{\gamma}$ are explicit affine functions of the exponents $\alpha_{i,k}$ in (22), as specified by (D.13) in Appendix D.
Theorem 5 provides a way for finding unknown exponents $\alpha_{i,k}$ in (22) (Theorem 3). The proof is shown in Appendix E.
Theorem 5.
Given Assumptions 1 to 6 and the consequent expression (22) of coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13), the vector
$\vec{\alpha}\coloneqq\begin{pmatrix}\vec{\alpha}_{1}\\ \vec{\alpha}_{2}\\ \vdots\\ \vec{\alpha}_{N_{\mathsf{l}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}},\quad\mbox{with}\quad\vec{\alpha}_{i}\coloneqq\begin{pmatrix}\alpha_{i,1}\\ \alpha_{i,2}\\ \vdots\\ \alpha_{i,N_{\mathsf{0}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{0}}},$ (25)
of exponents in (22) must satisfy the linear system
$\textbf{S}\,\vec{\alpha}=\vec{s},\quad\textbf{S}\in\mathbb{R}^{N_{\mathsf{r}}\times N_{\mathsf{c}}},\quad\vec{s}\in\mathbb{R}^{N_{\mathsf{r}}},\quad N_{\mathsf{c}}\coloneqq N_{\mathsf{l}}\,N_{\mathsf{0}},\quad N_{\mathsf{r}}\coloneqq N_{\mathsf{l}}\,N_{\mathsf{p}}-\mathrm{rank}(\textbf{Y}),$ (26)
where the entries of S and $\vec{s}$ are explicitly known real numbers identified by (E.5)-(E.6) in Appendix E. The matrix $\textbf{Y}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})}$ is explicitly known, as it is given in (24) and specified by (D.13) in Appendix D.
3.1.2 Optimization
Theorem 6 enables the computation of numerical values $\theta_{1},\dots,\theta_{N_{\mathsf{x}}}$ of the characteristic constants $\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}}$ that minimize the deviation $C_{U}$ from unity of the magnitudes of $\lambda_{i}$ (13), with $i\in U\subseteq\{1,\dots,N_{\mathsf{l}}\}$. The proof is shown in Appendix F.
Theorem 6.
Given Assumption 1 for coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13), $U\subseteq\{1,\dots,N_{\mathsf{l}}\}$ and any fixed choice of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$, the values of $(\theta_{1},\dots,\theta_{N_{\mathsf{x}}})\in(0,\infty)^{N_{\mathsf{x}}}$ attaining the global minimum of the function
$C_{U}(\theta_{1},\dots,\theta_{N_{\mathsf{x}}})\coloneqq\sum_{i\in U}\left(\log_{10}(\lambda_{i}(\tilde{\theta}_{1},\dots,\tilde{\theta}_{N_{\mathsf{x}}},\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}))\right)^{2},\quad C_{U}:(0,\infty)^{N_{\mathsf{x}}}\to[0,\infty),$ (27)
$\sum_{k=1}^{N_{\mathsf{x}}}\left(\sum_{i\in U}\,c_{i,k}\,c_{i,j}\right)\rho_{k}=-\sum_{i\in U}\,c_{i,j}\,\log_{10}(\kappa_{i}),\quad\forall j=1,\dots,N_{\mathsf{x}},$ (28)
where $\rho_{k}\coloneqq\log_{10}(\theta_{k})$ and the numerical value $\kappa_{i}\in(0,\infty)$ of each $\tilde{\kappa}_{i}=\tilde{\kappa}_{i}(\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}})$ is fixed by the choice of $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$.
Theorem 7 enables the computation of numerical values of the exponents $\alpha_{i,k}$ in (22) that minimize the deviation $C_{U}$ from unity of the magnitudes of $\lambda_{i}$, with $i\in U\subseteq\{1,\dots,N_{\mathsf{l}}\}$. The proof is shown in Appendix G.
Theorem 7.
Given any fixed choice of $\pi_{1},\dots,\pi_{N_{\mathsf{0}}}\in(0,\infty)$ and set $U\subseteq\{1,\dots,N_{\mathsf{l}}\}$, the values of $\vec{\alpha}\in\mathbb{R}^{N_{\mathsf{c}}}$ (25), $N_{\mathsf{c}}\coloneqq N_{\mathsf{l}}\,N_{\mathsf{0}}$, attaining the global minimum of the function
$C_{U}(\vec{\alpha})\coloneqq\sum_{i\in U}\left(\log_{10}(\lambda_{i})\right)^{2},\quad\lambda_{i}=\pi_{1}^{\alpha_{i,1}}\,\pi_{2}^{\alpha_{i,2}}\,\cdots\,\pi_{N_{\mathsf{0}}}^{\alpha_{i,N_{\mathsf{0}}}},\quad C_{U}:\mathbb{R}^{N_{\mathsf{c}}}\to[0,\infty),$ (29)
subject to a constraint of the form
$\boldsymbol{\Sigma}\,\vec{\alpha}=\vec{\sigma},\quad\boldsymbol{\Sigma}\in\mathbb{R}^{\hat{N_{\mathsf{r}}}\times N_{\mathsf{c}}},\quad\vec{\sigma}\in\mathbb{R}^{\hat{N_{\mathsf{r}}}},\quad\hat{N_{\mathsf{r}}}\in\mathbb{N}$, (30)
satisfy the system of equations
$\boldsymbol{\Sigma}\,\vec{\alpha}=\vec{\sigma},\quad\nabla C_{U}(\vec{\alpha})\,+\,\boldsymbol{\Sigma}^{T}\,\vec{\nu}=\vec{0},$ (31)
assuming that
$\mathrm{rank}(\boldsymbol{\Sigma})=\hat{N_{\mathsf{r}}}<N_{\mathsf{c}},$ (32)
and there is a vector $\vec{\nu}\in\mathbb{R}^{\hat{N_{\mathsf{r}}}}$ such that (31) holds.